OpenStudy (anonymous):

by using integration by part solve: {x(x+4)^-2 dx

OpenStudy (anonymous):

Let u = x and dv = 1/(x+4)^2 . Then, du = 1 and v = -1 / (x+4) Use integration by parts now.